2e^(5x+2) = 8
Thank you,
Alexus

Answer by stanbon(75887)

Divide both sides by "2" to get e^(5x+2)=4
Take the natural log of both sides to get:
5x+2=ln4
5x=1.39-2
5x=-0.61
x=-0.122
Cheers,
Stan H.